OpenStudy (anonymous):

medal&fan mhm A poster is shown below: A rectangle is shown. The length of the rectangle is labeled 12 feet. The width of the rectangle is labeled 3 feet. What are the dimensions if the poster is enlarged by a factor of five over two ? 7.5 ft by 30 ft 15 ft by 60 ft 1.5 ft by 6 ft 15 ft by 6 ft

OpenStudy (anonymous):

@StudyGurl14 maybe can ya help?

OpenStudy (studygurl14):

\(12\times\Large\frac{5}{2}\rightarrow\frac{60}{2}\rightarrow30\) by \(3\times\Large\frac{5}{2}\rightarrow\frac{15}{2}\rightarrow 7.5\)